Hi I am having trouble with this question. Can you help?
9^(4x-3)=4^(2x-4)

Answer by fractalier(2101)

This has to be solved using logarithms...so from
9^(4x-3) = 4^(2x-4)
take the log of both sides, apply the power rule and use your algebar tools to solve for x...here goes...
log 9^(4x-3) = log 4^(2x-4)
(4x-3)*log 9 = (2x-4)*log 4
4xlog9 - 3log9 = 2xlog4 - 4log4
4xlog9 - 2xlog4 = 3log9 - 4log4
x(4log9 - 2log4) = 3log9 - 4log4
x = (3log9 - 4log4) / (4log9 - 2log4)